# Tabular foundation model to detect empathy from visual cues

### v1

@@ 4.6. Limitations
We demonstrate the use of tabular foundation models on statistical attributes calculated from temporal video series. While effective, this approach may limit the model’s ability to capture patterns present in the raw temporal data. Incorporating the full temporal sequence into the tabular foundation model could potentially reveal additional patterns and further improve classification performance.
5. Conclusion
@@
This paper investigates the potential of two recent tabular foundation models for detecting empathy elicited through narrative storytelling. Given that raw video sequences are often unavailable due to privacy and ethical constraints, our approach focuses on learning from visual features. We evaluate both in-context learning (ICL) and fine-tuning approaches. Under the published evaluation protocol, both ICL and fine-tuning are competitive and outperform several strong baselines. In addition to the established protocol, we are the first to evaluate cross-subject generalisation in detecting empathy elicited through storytelling. Most models, including the ICL setup of the tabular foundation model, struggle in this setting. However, fine-tuning leads to significantly better generalisation across unseen subjects. Given the continued limitations in accessing raw video data due to privacy and ethical constraints, our results demonstrate practical values and future potential of tabular foundation models.
6. Acknowledgments

### v2

@@ IV-F Limitations
We demonstrate the use of TFMs on statistical attributes calculated from temporal video series. While effective, this approach may limit the model’s ability to capture patterns present in the raw temporal data. Incorporating the full temporal sequence into the TFM could potentially reveal additional patterns and further improve classification performance.
V Conclusion
@@
This paper proposes TFMPathy, a system utilising the potential of TFMs for detecting empathy elicited through narrative storytelling. Given that raw video sequences are often unavailable due to privacy and ethical constraints, TFMPathy learns from visual features and does not require any raw videos. Experimented on two recent TFMs (TabICL and TabPFN), we evaluate both ICL and fine-tuning approaches. In a public evaluation benchmark on human-robot interaction, TFMPathy in both ICL and fine-tuning settings is competitive and outperforms several strong baselines. In addition to the established protocol, we are the first to evaluate cross-subject generalisation in detecting empathy elicited through storytelling. While most models, including TFMPathy in the initial ICL setup, struggle in generalisation, TFMPathy in the fine-tuning setup leads to significantly better generalisation across unseen subjects. Given the persistent privacy and ethical constraints in sharing raw videos, our proposed TFMPathy system offers practical value and future potential for AI systems that rely on human-centred video datasets.
Acknowledgments